OpenStudy (anonymous):

Alicia and Wayne are both collecting coins. Alicia has three times as many coins as Wayne. Together they have 712 coins. How many coins does Wayne have?

OpenStudy (anonymous):

534

OpenStudy (anonymous):

no my bad. Alicia has 534 coins, Wayne has 178 coins

OpenStudy (anonymous):

how'd u get the answer?

OpenStudy (anonymous):

i can answer that

OpenStudy (anonymous):

its really easy take 712 divide it by 4 because Alicia has 3times more coins than wayne thats three and then add wayne in there thats 4 take 712/4=178 take 712-178=534 is how many alicia has see 178+178=356 then 356*2=712

OpenStudy (anonymous):

ohhhhh! thank you!!! i knew it was really simple, i just couldn't figure it out, anyways thanks!

OpenStudy (anonymous):